# Energy Matching: Unifying Flow Matching and Energy-Based Models for Generative Modeling

### v3

@@ 2.1.2 Contrastive objective $\operatorname{\mathcal{L}_{\mathrm{CD}}}$
2.1.3 Dual objective
To balance the flow-like loss $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ with the contrastive divergence loss $\operatorname{\mathcal{L}_{\mathrm{CD}}}$, we adopt the linear temperature schedule (3). We define a sampling time $\operatorname{\tau_{s}}$. In fact, although convergence to the equilibrium distribution is guaranteed only as $\operatorname{\tau_{s}}\to\infty$, we empirically observe that sample quality (measured with Fréchet inception distance (FID)) plateaus by $\operatorname{\tau_{s}}=3.0$ on CIFAR-10; see Section A.2. The time discretization determines the temperature samples $\varepsilon^{(m)}\vcentcolon=\varepsilon(m\Delta t)$. We define a dataset-specific hyperparameter $\lambda_{\mathrm{CD}}$ to scale $\operatorname{\mathcal{L}_{\mathrm{CD}}}$. The resulting algorithm is described in detail in Algorithm 1 and Algorithm 2. Since Algorithm 2 benefits from high-quality negatives, we begin with Algorithm 1 (and, thus, with $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ only) to ensure sufficient mixing of noise-initialized negatives. Section A.1 discusses how the landscape of $V_{\theta}$ evolves across these two phases. Hyper-parameters for each dataset are defined in Appendix D.
@@ 2.1.3 Dual objective
1:$\theta\leftarrow\theta_{\text{pretrained}}$triangleright Initialize from the warmed-up model (Algorithm 1)
2:for iteration $0,1,\ldots$ do
3:    Sample mini-batch $\{x_{\mathrm{data}}^{(j)}\}_{j=1}^{B}\sim\mathcal{D}$triangleright Mini-batch of data samples
4:    $\operatorname{\mathcal{L}_{\mathrm{OT}}}\leftarrow$ Use lines 4–7 from Algorithm 1
5:    Initialize $\{x_{\mathrm{neg}}^{(0)}\}$ from noise and/or data $\triangleright$ Negative samples
6:    for $m=0,1,\ldots,M_{\text{Langevin}}$ do
7:         Sample $\eta\sim\mathcal{N}(0,I)$
8:         $x_{\mathrm{neg}}^{(m+1)}\leftarrow x_{\mathrm{neg}}^{(m)}-\Delta t\nabla_{x}V_{\mathrm{sg}(\theta)}(x_{\mathrm{neg}}^{(m)})+\sqrt{2\Delta t\varepsilon^{(m)}}\eta$triangleright Langevin dynamics given $\varepsilon(t)$
9:         if $x_{\mathrm{neg}}^{(0)}$ initialized from data then
10:             Use $\varepsilon_{\max}$ instead of $\varepsilon^{(m)}$triangleright Use maximum temperature if initialized at the data manifold
11:         end if
12:    end for
13:    $\mathcal{L}_{\mathrm{neg}}\leftarrow\mathbb{E}[V_{\theta}(x_{\mathrm{neg}}^{(M)})]$
14:    $\mathcal{L}_{\mathrm{pos}}\leftarrow\mathbb{E}_{\mathrm{data}}[V_{\theta}(x_{\mathrm{data}})]$$\triangleright$ Compute expected potential energy on the data
15:    $\operatorname{\mathcal{L}_{\mathrm{CD}}}\leftarrow\mathcal{L}_{\mathrm{pos}}-\mathcal{L}_{\mathrm{neg}}$
16:    $\mathcal{L}(\theta)\leftarrow\operatorname{\mathcal{L}_{\mathrm{OT}}}+\lambda_{\mathrm{CD}}\operatorname{\mathcal{L}_{\mathrm{CD}}}$
17:    Update $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}(\theta)$triangleright Gradient descent step
18:end for
19:return $\theta$triangleright Trained $\theta$
@@ 3 Applications
In this section, we demonstrate the effectiveness and versatility of our proposed Energy Matching approach across three applications: (i) unconditional generation (ii) inverse problems, and (iii) LID estimation. The model architecture and all the training details are reported in Appendix D.
@@ 3.1 Unconditional generation
We compare four classes of generative models: (1) Diffusion models, which deliver state-of-the-art quality but typically require many sampling steps; (2) Flow-based methods, which learn OT paths for more efficient sampling with fewer steps; (3) EBMs, which directly model the log-density as a scalar field, offering flexibility for inverse problems and constraints but sometimes at the expense of sample quality; and (4) Ensembles (Diffusion with one or many EBMs), which combine diffusion’s robust sampling with elements of EBM flexibility but can become large and complex to train. Our approach, Energy Matching, offers a simple (a single time-independent scalar field) yet powerful EBM-based framework. We evaluate our approach on CIFAR-10 (Krizhevsky et al., 2009) and ImageNet32x32 (Deng et al., 2009; Chrabaszcz et al., 2017) datasets, reporting FID scores in Table 1 and Table 2. Our method surpasses state-of-the-art EBMs, achieving a $50\%$ reduction in FID.
@@
. Learning Unnormalized Data Likelihood Learning Transport/Score Along Noised Trajectories Ensembles: Diffusion + (one or many) EBMs Diffusion Models Hierarchical EBM Diffusion (Cui and Han, 2024) 8.93 DDPM (Ho et al., 2020) 6.54 EGC (Guo et al., 2023) 5.36 DDPM++ (62M params, 1000 steps) (Kim et al., 2021) 3.45 Cooperative DRL (40M params) (Zhu et al., 2024) 4.31 NCSN++ (107M params, 1000 steps) (Song et al., 2021) 2.45 Cooperative DRL-large (145M params) (Zhu et al., 2024) 3.68 Energy-based Models Flow-based Models ImprovedCD (Du et al., 2021) 25.1 Action Matching (Neklyudov et al., 2023) 10.07 CLEL-large (32M params) (Lee et al., 2023) 8.61 Flow-matching (Lipman et al., 2023) 6.35 Energy Matching (50M params, Ours) 3.97 OT-CFM (37M params) (Tong et al., 2023) 4.11
@@
Energy Matching (50M params, Ours) 7.67
@@ 3.2 Inverse problems
where $E_{\theta}(x)$ is an energy function which one can learn from the data, an EBM. Because we want to sample $x$ given a measurement $y$, this reconstruction task is often referred to as an inverse problem. Here, $\|y-Ax\|^{2}$ encodes the measurement fidelity with $\zeta$ controlling the balance between this fidelity term and the prior. We obtain the prior term $E_{\theta}(x)=\frac{V_{\theta}(x)}{\varepsilon_{\max}}$ by training $V_{\theta}(x)$ via Energy Matching. Samples from this posterior can be drawn by starting from a random sample $x^{(0)}\sim\mathcal{N}(0,I)$ and following a Langevin update. We detail the algorithm for generating solutions to inverse problems in Algorithm 3 of Appendix B. We demonstrate our model’s capabilities qualitatively through a controlled inpainting task and quantitatively via a protein inverse design benchmark.
@@
Suppose we want to recover two images from a masked image while encouraging diverse reconstructions. EBMs allow this by introducing an additional interaction energy, $W(x_{1},x_{2})=-\frac{\|B(x_{1}-x_{2})\|^{2}}{\sigma^{2}}$, where $B$ has ones in the region of interest (focusing diversity there) and zeros elsewhere, and $\sigma$ is a hyper-parameter controlling the interaction’s strength. Specifically, we define $p(x_{1},x_{2}\mid y)\propto p(x_{1}\mid y)\,p(x_{2}\mid y)\,\exp\bigl(-W(x_{1},x_{2})\bigr{missing})$, which gives high probability to pairs $(x_{1},x_{2})$ that lie far apart in the specified region $B$. This encourages exploring the edges of the posterior rather than just its modes, and with suitable $W$, samples shift toward rare events without needing many draws. To illustrate the interaction term’s advantages for diverse reconstruction, we apply our method to a CelebA (Liu et al., 2015) $64\times 64$ inpainting task. As shown in Figure 2, we start from a partially observed (masked) face and aim to reconstruct two distinct high-fidelity completions.
@@
Figure 2: Controlled inpainting for diverse reconstructions. On the left is the masked face. On the right are two reconstructions: the top pair without the interaction term and the bottom pair with it. The interaction term applies in the solid red square (where $B$ has ones), and the measurement matrix $A$ is the dotted blue square (zeros inside, ones outside). By encouraging $x_{1}$ and $x_{2}$ to differ in the target region, the interaction yields a wider range of completions while preserving fidelity.
@@
In Figure 3, we demonstrate our method’s performance on the inverse design problem of generating Adeno-Associated Virus (AAV) capsid protein segments (Bryant et al., 2021). Given a desired functional property (fitness)—here defined as the predicted viral packaging efficiency normalized between 0 and 1—the goal is to design novel protein sequences satisfying this target condition. Beyond achieving high fitness, practical inverse design requires generating diverse candidate sequences to ensure robustness in the downstream experimental validation (Jain et al., 2022). We evaluate on two benchmark splits (medium and hard), which correspond to subsets of the original AAV dataset differing in baseline fitness distributions and required mutational distance from known high-performing variants (Kirjner et al., 2024). Leveraging the latent-space representation of VLGPO (Bogensperger et al., 2025), we employ our Energy-Matching Langevin sampler with an inference-time tunable repulsion term, allowing explicit control over the diversity of the designed proteins. This enables a flexible trade-off between fitness and diversity, resulting in high fitness scores alongside substantially improved sequence diversity. See Section B.2 for experimental details and dataset descriptions.
@@ 3.3 Local intrinsic dimension estimation
Real-world datasets, despite displaying a high number of variables, can often be represented by lower-dimensional manifolds—a concept referred to as the manifold hypothesis (Fefferman et al., 2016). The dimension of such a manifold is called the intrinsic dimension. Estimating the LID at a given point reveals its effective degrees of freedom or directions of variation, offering insight into data complexity and adversarial vulnerabilities. We defer the precise definition to Appendix C.
@@
Recent work leverages pre-trained diffusion models to estimate the LID (Kamkari et al., 2024; Stanczuk et al., 2024) by examining the learned score function. However, because these models do not learn the score at $t=0$ (the data manifold), their estimates become unreliable there. As a result, current methods rely on approximations, for instance, by evaluating the score at small but nonzero times $t=t_{0}$, where the computations remain sufficiently reliable.
@@
Unlike diffusion models, EBMs explicitly parametrize the relative data likelihood. This explicit parametrization enables efficient analysis of the curvature of the underlying data manifold – in this example, estimating the LID. To this end, we compute the Hessian matrix $\nabla_{x}^{2}V(x_{\mathrm{data}})$ at a given data point and perform its spectral decomposition. We define near-zero eigenvalues as those whose absolute values lie within a small threshold $\tau$ (in our experiments, we set $\tau=10$ for MNIST (Deng, 2012) and $\tau=2$ for CIFAR-10). The count of near-zero eigenvalues reflects the number of flat directions and thus reveals the local dimension. As shown in Table 3, the LID estimates we obtain exhibit stronger correlations with PNG compression size [^2] PNG is a lossless compression scheme specialized for images and can provide useful guidance when no LID ground truth is available Kamkari et al. (2024). (evaluated on 4096 images) using Spearman’s correlation. Figure 4 offers qualitative illustrations. Our EBM-based approach compares favorably to diffusion-based methods, as it relies on fewer approximations by performing computations exactly on the data manifold rather than merely in its vicinity.
@@
Energy Matching (Ours) 0.895 0.901
@@ 4 Conclusion and limitations
First, our method requires the extra gradient computation with respect to input, which can increase GPU memory usage (e.g. 20-40%), in particular during training. Second, for very high-resolution datasets, computing the full Hessian spectrum ($O(N^{3})$) may be impractical; and partial spectrum methods like random projections or iterative solvers may be used instead.
@@
In contrast to the widespread belief, we demonstrated that time-independent curl-free methods for generative flows work and offer an exciting direction for future research. Our Energy Matching approach has the potential to yield novel insights into controlled generation and inverse problems for cancer research Weidner et al. (2024); Balcerak et al. (2024), molecules and proteins Wu et al. (2022); Bilodeau et al. (2022), computational fluid dynamics Gao et al. (2024); Shysheya et al. (2024); Molinaro et al. (2024) and other fields where precise control over generated samples and effective integration of priors or constraints are crucial.

### v5

@@ 2.1.2 Contrastive objective $\operatorname{\mathcal{L}_{\mathrm{CD}}}$
2.1.3 Dual objective and implementation notes
To balance the deterministic flow-like regime (where $\varepsilon\approx 0$) away from the data manifold and the stochastic Boltzmann regime (where $\varepsilon\approx\varepsilon_{\max}$) near equilibrium, we adopt the linear temperature schedule described in (3). We introduce a dataset-specific hyperparameter $\lambda_{\mathrm{CD}}$ to stabilize the contrastive objective by appropriately weighting $\operatorname{\mathcal{L}_{\mathrm{CD}}}$ relative to $\operatorname{\mathcal{L}_{\mathrm{OT}}}$. The resulting algorithm is described in detail in Algorithm˜1 and Algorithm˜2. Since Algorithm˜2 benefits from high-quality negatives, we begin with Algorithm˜1 (and, thus, with $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ only) to ensure sufficient mixing of noise-initialized negatives.
Given the trained models, we define a sampling time $\operatorname{\tau_{s}}$. Although convergence to the equilibrium distribution is guaranteed only as $\operatorname{\tau_{s}}\to\infty$, we empirically observe that sample quality (measured with Fréchet inception distance (FID)) plateaus by $\operatorname{\tau_{s}}=3.25$ on CIFAR-10; see Section˜A.2. The sampling procedure, which optionally includes conditional and interaction terms, is detailed in Algorithm˜3. In practice, we implement training using explicit Euler–Maruyama updates and sampling with an Euler–Heun predictor-corrector scheme, while for simplicity the algorithms illustrate only explicit updates. Additionally, the constant factor $1/\varepsilon_{\max}$ in $\operatorname{\mathcal{L}_{\mathrm{CD}}}$ is absorbed into $\lambda_{\mathrm{CD}}$.
Section˜A.1 discusses how the landscape of $V_{\theta}$ evolves across these two phases. Hyperparameters for each dataset, along with intuitions to guide their selection for new datasets, are provided in Appendix˜D.
@@ 2.1.3 Dual objective and implementation notes
[…]
2:for iteration $n=0,1,\ldots$ do
3:  $\operatorname{\mathcal{L}_{\mathrm{OT}}}\leftarrow$ Use lines 3–8 from Algorithm 1
4:  Initialize negative samples $\{x_{\mathrm{neg},b}^{(0)}\}_{b=1}^{B}$ from noise and/or data $\triangleright$ Negative samples
5:  for $m=0,1,\ldots,M_{\text{Langevin}}-1$ do
6:    for $b=1$ to $B$ do
[…]
8:     Sample $\eta_{b}\sim\mathcal{N}(0,I)$
9:     $x_{\mathrm{neg},b}^{(m+1)}\leftarrow x_{\mathrm{neg},b}^{(m)}-\Delta t\nabla_{x}V_{\mathrm{sg}(\theta)}(x_{\mathrm{neg},b}^{(m)})+\sqrt{2\Delta t\varepsilon^{(m)}}\,\eta_{b}$ $\triangleright$ Langevin dynamics step
10:    end for
11:  end for
12:  $\operatorname{\mathcal{L}_{\mathrm{CD}}}\leftarrow\frac{1}{B}\sum_{b=1}^{B}\left[V_{\theta}(x_{\mathrm{data},b})-V_{\theta}(x_{\mathrm{neg},b}^{(M_{\text{Langevin}})})\right]$ $\triangleright$ Contrastive divergence loss
13:  $\mathcal{L}(\theta)\leftarrow\operatorname{\mathcal{L}_{\mathrm{OT}}}+\lambda_{\mathrm{CD}}\operatorname{\mathcal{L}_{\mathrm{CD}}}$
14:  Update $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}(\theta)$ $\triangleright$ Gradient descent step
15:end for
16:return $\theta$ $\triangleright$ Trained $\theta$
@@ 3 Applications
In this section, we demonstrate the effectiveness and versatility of our proposed Energy Matching approach across three applications: (i) unconditional generation (ii) inverse problems, and (iii) LID estimation. The model architecture and all the training details are reported in Appendix˜D.
@@ 3.1 Unconditional generation
We compare four classes of generative models: (1) Diffusion models, which deliver state-of-the-art quality but typically require many sampling steps; (2) Flow-based methods, which learn OT paths for more efficient sampling with fewer steps; (3) EBMs, which directly model the log-density as a scalar field, offering flexibility for inverse problems and constraints but sometimes at the expense of sample quality; and (4) Ensembles (Diffusion with one or many EBMs), which combine diffusion’s robust sampling with elements of EBM flexibility but can become large and complex to train. Our approach, Energy Matching, offers a simple (a single time-independent scalar field) yet powerful EBM-based framework. We evaluate our approach on CIFAR-10 (Krizhevsky and Hinton, 2009) and ImageNet32x32 (Deng et al., 2009; Chrabaszcz et al., 2017) datasets, reporting FID scores in Table˜1 and Table˜2, respectively. Our method outperforms state-of-the-art EBMs, reducing the FID score by more than $50\%$.
@@
. Learning Unnormalized Data Likelihood Learning Transport/Score Along Noised Trajectories Ensembles: Diffusion + (one or many) EBMs Diffusion Models Hierarchical EBM Diffusion (Cui and Han, 2024) 8.93 DDPM (Ho et al., 2020) 6.45 EGC (Guo et al., 2023) 5.36 DDPM++ (62M params, 1000 steps) (Kim et al., 2021) 3.45 Cooperative DRL (40M params) (Zhu et al., 2024) 4.31 NCSN++ (107M params, 1000 steps) (Song et al., 2021) 2.45 Cooperative DRL-large (145M params) (Zhu et al., 2024) 3.68 Energy-based Models Flow-based Models ImprovedCD (Du et al., 2021) 25.1 Action Matching (Neklyudov et al., 2023) 10.07 CLEL-large (32M params) (Lee et al., 2023) 8.61 Flow-matching (Lipman et al., 2023) 6.35 Energy Matching (50M params, Ours) 3.34 OT-CFM (37M params) (Tong et al., 2023) 4.04
@@
Energy Matching (50M params, Ours) 6.64
@@ 3.2 Inverse problems
where $E_{\theta}(x)$ is an energy function which one can learn from the data, an EBM. Because we want to sample $x$ given a measurement $y$, this reconstruction task is often referred to as an inverse problem. Here, $\|y-Ax\|^{2}$ encodes the measurement fidelity with $\zeta$ controlling the balance between this fidelity term and the prior. We obtain the prior term $E_{\theta}(x)=\frac{V_{\theta}(x)}{\varepsilon_{\max}}$ by training $V_{\theta}(x)$ via Energy Matching. Samples from this posterior can be drawn by starting from a random sample $x^{(0)}\sim\mathcal{N}(0,I)$ and following a Langevin update. We detail the algorithm for generating solutions to inverse problems in Algorithm˜3 (which also incorporates additional interaction energy $W(x,x^{\prime})$ between generated samples). We demonstrate our model’s capabilities qualitatively through a controlled inpainting task and quantitatively via a protein inverse design benchmark. Specific hyperparameters are detailed in Appendix˜D.
Algorithm 3 Unconditional/conditional sampling with optional interaction energy
1:for $m=1$ to $M$ do
2:  Initialize $x_{m}^{(0)}$ from noise and/or data $\triangleright$ Initialize each chain
3:end for
[…]
5:for $n=0,1,\ldots,N-1$ do
6:  for $m=1,2,\ldots,M$ do $\triangleright$ Prior + data fidelity + interaction
[…]
8:    $U_{\theta}\bigl(x_{m}^{(n)}\bigr)\leftarrow V_{\theta}\bigl(x_{m}^{(n)}\bigr)+\varepsilon^{(n)}\bigl\lVert y-A\bigl(x_{m}^{(n)}\bigr)\bigr\rVert^{2}/\zeta^{2}+\varepsilon^{(n)}\sum_{k\neq m}W\bigl(x_{m}^{(n)},x_{k}^{(n)}\bigr)$
9:    Sample $\eta_{m}^{(n)}\sim\mathcal{N}(0,I)$
10:    $x_{m}^{(n+1)}\leftarrow x_{m}^{(n)}-\Delta t\,\nabla_{x}U_{\theta}\bigl(x_{m}^{(n)}\bigr)+\sqrt{2\varepsilon^{(n)}\Delta t}\;\eta_{m}^{(n)}$ $\triangleright$ Langevin dynamics step
11:  end for
12:end for
13:return $\{x_{m}^{(N)}\}_{m=1}^{M}$ $\triangleright$ Final samples
@@
[…]
Figure 2: Controlled inpainting for diverse reconstructions. On the left is the masked face. On the right are two reconstructions: the top pair without the interaction term and the bottom pair with it. The interaction term applies in the solid red square (where $B$ has ones), and the measurement matrix AA is the dotted blue square (zeros inside, ones outside). By encouraging $x_{1}$ and $x_{2}$ to differ in the target region, the interaction yields a wider range of completions while preserving fidelity.
@@
In Figure˜3, we demonstrate our method’s performance on the inverse design problem of generating Adeno-Associated Virus (AAV) capsid protein segments (Bryant et al., 2021). Given a desired functional property (fitness)—here defined as the predicted viral packaging efficiency normalized between 0 and 1—the goal is to design novel protein sequences satisfying this target condition. Beyond achieving high fitness, practical inverse design requires generating diverse candidate sequences to ensure robustness in the downstream experimental validation (Jain et al., 2022). We evaluate on two benchmark splits (medium and hard), which correspond to subsets of the original AAV dataset differing in baseline fitness distributions and required mutational distance from known high-performing variants (Kirjner et al., 2024). Leveraging the latent-space representation of VLGPO (Bogensperger et al., 2025), we employ our Energy-Matching Langevin sampler with an inference-time tunable repulsion term, allowing explicit control over the diversity of the designed proteins. This enables a flexible trade-off between fitness and diversity, resulting in high fitness scores alongside substantially improved sequence diversity. See Appendix˜B for experimental details and dataset descriptions.
@@ 3.3 Local intrinsic dimension estimation
Real-world datasets, despite displaying a high number of variables, can often be represented by lower-dimensional manifolds—a concept referred to as the manifold hypothesis (Fefferman et al., 2016). The dimension of such a manifold is called the intrinsic dimension. Estimating the LID at a given point reveals its effective degrees of freedom or directions of variation, offering insight into data complexity and adversarial vulnerabilities. We defer the precise definition to Appendix˜C.
@@
Recent work leverages pretrained diffusion models to estimate the LID (Kamkari et al., 2024; Stanczuk et al., 2024) by examining the learned score function. However, since these models do not learn the score at the data manifold ($t=1$), their estimates become unreliable there. Consequently, current methods rely on approximations, for instance by evaluating the score in the proximity of the data manifold ($t=1-t_{0}$), where computations remain sufficiently reliable.
@@
Unlike diffusion models, EBMs explicitly parametrize the relative data likelihood. This explicit parametrization enables efficient analysis of the curvature of the underlying data manifold – in this example, estimating the LID. To this end, we compute the Hessian matrix $\nabla_{x}^{2}V(x_{\mathrm{data}})$ at a given data point and perform its spectral decomposition. We define near-zero eigenvalues as those whose absolute values lie within a small threshold $\tau$ (in our experiments, we set $\tau=3$ for MNIST (Deng, 2012) and $\tau=2$ for CIFAR-10). The count of near-zero eigenvalues reflects the number of flat directions and thus reveals the local dimension. As shown in Table˜3, the LID estimates we obtain exhibit stronger correlations with PNG compression size [^2] PNG is a lossless compression scheme specialized for images and can provide useful guidance when no LID ground truth is available Kamkari et al. (2024). (evaluated on 4096 images) using Spearman’s correlation. Figure˜4 offers qualitative illustrations. Our EBM-based approach compares favorably to diffusion-based methods, as it relies on fewer approximations by performing computations exactly on the data manifold rather than merely in its vicinity.
@@
Energy Matching (Ours) 0.877 0.901
@@ 4 Conclusion and limitations
First, our method requires an additional gradient computation with respect to the input, which can increase GPU memory usage (e.g., by 20–40%), particularly during training. Second, when estimating the LID (Section 3.3) for very high-dimensional datasets, computing the full Hessian spectrum may be impractical due to its computational complexity of $O(d^{3})$; in such cases, partial-spectrum methods such as random projections or iterative solvers can be employed instead.
@@
Contrary to widespread belief, we demonstrated that time-independent irrotational methods for generative flows are highly effective and offer an exciting direction for future research. Our Energy Matching approach has the potential to yield novel insights into controlled generation and inverse problems for cancer research Weidner et al. (2024); Balcerak et al. (2024), molecules and proteins Wu et al. (2022); Bilodeau et al. (2022), computational fluid dynamics Gao et al. (2024); Shysheya et al. (2024); Molinaro et al. (2024), and other fields where precise control over generated samples and effective integration of priors or constraints are crucial. Moreover, Energy Matching aligns naturally with recent generative AI trends toward scaling inference for new capabilities (Zhang et al., 2025; Ma et al., 2025), further broadening its potential impact across scientific and engineering domains.
Acknowledgments and Disclosure of Funding
This research was supported by the Helmut Horten Foundation and the European Cooperation in Science and Technology (COST).